# Corks, exotic 4-manifolds and genus functions

### v3

@@ 10. Exotically knotted submanifolds
We identify the boundary $\partial E(X)$ with $M$ using a fixed diffeomorphism $Y\to E(X)$. Note that $Y$ is oriented due to our convention. Let $\mathbb{P}$ be the invertible homology cobordism from $M^{\prime}$ to some closed (connected) 3-manifold $N$ used in Theorem 3.6. We denote the 4-manifold $C\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}$ by $S_{\mathrm{id}_{M^{\prime}}}$. We attach the upside down cobordism $\overline{\mathbb{P}}$ to $S_{\mathrm{id}_{M^{\prime}}}$ along $N$ using $\mathrm{id}_{N}$, and then attach the upside down cobordism $\overline{P}$ of $P$ along $M^{\prime}$ using $\mathrm{id}_{M^{\prime}}$. We denote the resulting 4-manifold $Y\cup_{\mathrm{id}_{M}}P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$. Since the double $\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}$ admits a diffeomorphism to $I\times M^{\prime}$ that fixes the boundary $\overline{M^{\prime}}\sqcup M^{\prime}$ pointwise (see Remark 3.9), the submanifold $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ admits a diffeomorphism to the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the boundary $\overline{M}\sqcup M$ pointwise. Let $\widehat{D}_{P}$ denote the preimage of $D_{P}$ under this diffeomorphism. By Lemma 10.3, the 4-manifold $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ is diffeomorphic to $Y$ and hence to the exterior $E(X)$, and this diffeomorphism $\varphi_{\mathrm{id}_{M^{\prime}}}:\widehat{Y}_{\mathrm{id}_{M^{\prime}}}\to E(X)$ fixes the boundary $M$ pointwise. We set $D_{P}^{0}=\varphi_{\mathrm{id}_{M^{\prime}}}(\widehat{D}_{P})$.
For each $f\in A$, we apply generalized cork twists to $S_{\mathrm{id}_{M^{\prime}}}$ and $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ along $(C,f)$, and denote the resulting 4-manifolds $C\cup_{f}\mathbb{P}$ and $C\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $S_{f}$ and $\widehat{Y}_{f}$, respectively. By the diffeomorphism $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ in the last paragraph, each cobordism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ is obtained from the exterior of $\widehat{D}_{P}$ in $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by attaching a homotopy 4-ball diffeomorphic to the partial twisted double of $C_{0}$. Thus, by Lemma 10.3, we have a diffeomorphism $P\cup_{f}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}\to P\cup_{{\mathrm{id}_{M^{\prime}}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the exterior of $\widehat{D}_{P}$ and hence the boundary $\overline{M}\sqcup M$ pointwise. Extending this diffeomorphism trivially, we obtain a diffeomorphism $\psi_{f}:\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that is identity on $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$. Let $\varphi_{f}:\widehat{Y}_{f}\to E(X)$ denote the diffeomorphism defined by $\varphi_{f}=\varphi_{\mathrm{id}_{M^{\prime}}}\circ\psi_{f}$.
For each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$, we denote the submanifolds $\varphi_{f}(S_{f})$, $\varphi_{f}(\overline{\mathbb{P}})$ and $\varphi_{f}(\overline{P})$ of $E(X)$ respectively by $S_{f}^{0}$, ¯$\overline{\mathbb{P}}^{0}_{f}$ and $\overline{P}^{0}_{f}$. Let $X_{f}$ be the submanifold ¯$\overline{\mathbb{P}}^{0}_{f}\cup\overline{P}^{0}_{f}\cup X$ of $Z$. Then, $X_{f}$’s are pairwise diffeomorphic since each $\varphi_{f}$ fixes the boundary $M$ pointwise. Note that they are pairwise orientedly diffeomorphic in the case where $Z$ is orientable. Clearly, the submanifold $\overline{\mathbb{P}^{0}_{f}}\cup\overline{P}^{0}_{f}$ of $Z$ is an invertible homology cobordism from $\overline{M}$ to $\overline{N}$, and each $X_{f}$ is obtained from $X$ by attaching this cobordism along $\overline{M}$. We thus see that each $X_{f}$ deformation retracts to $X$ and topologically resembles $X$ by Lemmas 3.8 and 10.3.
We will show that the pairs $X_{f}$’s are pairwise exotically knotted and exotically embedded. We note that $Z=S_{f}^{0}\cup X_{f}$ for each $f\in A\sqcup\{{\mathrm{id}_{M^{\prime}}}\}$. By Theorem 3.6, the exteriors of $X_{f}$’s are pairwise exotic, and thus the pairs $(Z,X_{f})$’s are pairwise non-diffeomorphic. Consequently, $X_{f}$’s are pairwise smoothly non-isotopic.
@@
In the rest, we will prove that $X_{f}$’s are pairwise topologically ambient isotopic by showing that each $X_{f}$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. Due to the assumption, each $f^{-1}\in A$ extends to a self-homeomorphism of $C$ that is identity on the exterior of the contractible 4-manifold $C_{0}$. Extending this homeomorphism trivially, we get a homeomorphism $\widehat{Y}_{f}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that maps $S_{f}$ to $S_{\mathrm{id}_{M^{\prime}}}$. Note that this homeomorphism is identity on the exterior $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$. Composing this homeomorphism with the diffeomorphism $\psi_{f}$, we obtain a self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that fixes the $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{P}$ and hence the boundary $M$ pointwise. Therefore, by composing this self-homeomorphism of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ with the diffeomorphisms $\varphi_{\mathrm{id}_{M^{\prime}}}$ and $\varphi_{\mathrm{id}_{M^{\prime}}}^{-1}$, we obtain a self-homeomorphism of $E(X)$ that maps $S_{f}^{0}$ to $S_{{\mathrm{id}_{M^{\prime}}}}^{0}$ and fixes $E(X)-\textnormal{int}\,D_{P}^{0}$ and the boundary $M$ pointwise. Extending this self-homeomorphism trivially, we obtain a self-homeomorphism ${\Phi}_{f}$ of $Z$ that maps $X_{f}$ to $X_{{\mathrm{id}_{M^{\prime}}}}$ and fixes the exterior of $D_{P}^{0}$ pointwise. Note that ${\Phi}_{f}$ fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
@@
The restriction of ${\Phi}_{f}$ to $D_{P}^{0}$ fixes the boundary pointwise. Since $D_{P}^{0}$ is diffeomorphic to $D^{4}$, it thus follows from a result of Orson and Powell [55, Corollary D] that the restriction to $D_{P}^{0}$ is topologically isotopic to the identity map on $D_{P}^{0}$ fixing the boundary pointwise. Thus, ${\Phi}_{f}$ is topologically isotopic to the identity map on $Z$. Note that this isotopy fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty. Therefore, each submanifold $X_{f}$ of $Z$ is topologically isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. This completes the proof. ∎
@@
Let $Y$ be a 4-manifold diffeomorphic to the exterior $E(X)$. We first consider the case where $Y$ is diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Let $(C_{0},\tau)$ be an infinite order contractible cork of Gompf (see Subsection 3.5). By results of Akbulut [5], Tange [64] and Naylor and Schwartz [54], the twisted double $C_{0}\cup_{\tau^{i}}\overline{C_{0}}$ is diffeomorphic to $S^{4}$ for each integer $i$. Therefore, we can apply Proposition 10.4 using a generalized cork obtained from $Y$ by Proposition 4.2. The claim thus follows in this case.
In the case where $Y$ is a 2-handlebody, we first modify $Y$ into a Stein handlebody $Y^{\prime}$ by using Proposition 3.3. Note that $Y$ is obtained from $Y^{\prime}$ by attaching the upside down cobordism of an invertible cobordism. Furthermore, $\partial Y$ is a deformation retract of this invertible homology cobordism. We attach this invertible cobordism to $X$ and denote the resulting 4-manifold by $X^{\prime}$. We can regard $X^{\prime}$ as a submanifold of $Z$ containing $X$ as a submanifold. Clearly, $X$ is a deformation retract of $X^{\prime}$. Therefore, we can prove the claim by applying the argument of the first case to the submanifold $X^{\prime}$. ∎
@@
As shown in the proof, each $X_{i}$ in Theorem 10.2 is obtained from $X$ by attaching an invertible homology cobordism, and thus the genus function of $X_{i}$ is algebraically equivalent to that of $X$.
@@
Let $M_{1},M_{2}$ be codimension-1 submanifolds of a (possibly non-orientable) 4-manifold $Z$. The manifold $Z$ is possibly non-orientable and may have ((((possibly disconnected)))) boundary. The submanifold $M_{1}$ will be called homology cobordant to $M_{2}$ through $Z$ if $Z$ admits a codimension-00 submanifold $H$ with $\partial H=\overline{M_{1}}\sqcup M_{2}$ such that $H$ is a homology cobordism from $M_{1}$ to $M_{2}$.
As an application of Theorem 10.2, we modify separating codimension-$1$ submanifolds into infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds. The theorem below is a refinement of Theorem 1.14.
@@
Let $Z$ be a ((((possibly non-orientable)))) $4$-manifold which may have ((((possibly disconnected)))) boundary, and let $M$ be a closed $3$-manifold embedded in the interior of $Z$ as a separating submanifold. Suppose the exterior of the tubular neighborhood of $M$ has a connected component which is either a ((((orientable)))) $2$-handlebody or diffeomorphic to a submanifold of a symplectic $4$-manifold with weakly convex boundary. Then $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$ through $Z$.
@@
Due to the assumption, we can apply Theorem 10.2 to a connected component $X$ of the exterior of the tubular neighborhood for $M$. Let $X_{i}$’s be the resulting infinitely many pairwise exotically knotted and exotically embedded codimension-0 submanifolds of $Z$, and let each $M_{i}$ be the boundary $\partial E(X_{i})$ of the exterior $E(X_{i})$. By the proof of Theorem 10.2, each $M_{i}$ is homology cobordant to $M$ through $Z$. Since $X_{i}$’s are pairwise topologically ambient isotopic, their boundary components $M_{i}$’s are also pairwise topologically ambient isotopic. By the proof of Theorem 10.2, the exteriors $E(X_{i})$’s are pairwise non-diffeomorphic, and $X_{i}$’s are pairwise diffeomorphic. Note that the exterior of the tubular neighborhood of each $M_{i}$ is the disconnected 4-manifold $E(X_{i})\sqcup X_{i}$. If a pair $(Z,M_{i})$ is diffeomorphic to some $(Z,M_{j})$, then $E(X_{i})\sqcup X_{i}$ is diffeomorphic to $E(X_{j})\sqcup X_{j}$ due to the uniqueness of the tubular neighborhood. However, in the case where $E(X_{i})$ is not homeomorphic to $X_{i}$, it follows that all $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic, and thus the pairs $(Z,M_{i})$’s are pairwise non-diffeomorphic. Therefore, $M_{i}$’s are pairwise exotically embedded and exotically knotted. In the case where $E(X_{i})$ is homeomorphic to $X_{i}$, at least infinitely many of $(E(X_{i})\sqcup X_{i})$’s are pairwise non-diffeomorphic by the pigeonhole principle. Therefore, at least infinitely many of $M_{i}$’s are pairwise exotically embedded and exotically knotted. ∎
The corollaries below are restatements of Corollaries 1.15 and 1.16.
@@
Suppose a closed $3$-manifold admits an embedding into $S^{4}$. Then every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds homology cobordant to the $3$-manifold.
@@
Let $M$ be a closed 3-manifold embedded in $S^{4}$. It follows from the Mayer–Vietoris exact sequence that $M$ is separating. Let $Y$ be a connected component of the exterior of the tubular neighborhood for $M$. Since $M$ is separating, the boundary $\partial Y$ is diffeomorphic to $M$. Also, $Y$ admits an embedding into $D^{4}$ and hence into any symplectic 4-manifold with weakly convex boundary. Now let $Z$ be an arbitrary 4-manifold. Then $Y$ admits an embedding into the interior of $Z$. Let $M$ denote the boundary of $Y$. Then we can apply Theorem 10.7 to this $M$, and thus the claim follows. ∎
@@
There exist infinitely many pairwise not homology cobordant closed $3$-manifolds such that, for each $3$-manifold of them, every ((((possibly non-orientable)))) $4$-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to the $3$-manifold.
@@
There are infinitely many closed 3-manifolds admitting embeddings into $S^{4}$ such that their homology cobordism classes are pairwise distinct. For example, $n(S^{1}\times S^{2})$ admits an embedding into $S^{4}$ for each $n\geq 0$. The claim thus follows from Corollary 10.8. ∎
The corollary below is a refinement of Corollary 1.18. Note that every closed 3-manifold embedded in a simply connected closed 4-manifold is separating as easily seen from the Mayer–Vietoris exact sequence.
@@
For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any closed $4$-manifold with $b_{2}<n$ as a separating submanifold.

### v5

@@ 10. Exotically knotted submanifolds
We identify the boundary $\partial E(X)$ with $M$ using a fixed diffeomorphism $Y\to E(X)$. Note that $Y$ is oriented due to our convention. Let $\mathbb{P}$ be the invertible homology cobordism from $M^{\prime}$ to some closed (connected) 3-manifold $N$ used in Theorem 3.9. We denote the 4-manifold $C\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}$ by $S_{\mathrm{id}_{M^{\prime}}}$. We attach the upside down cobordism $\overline{\mathbb{P}}$ to $S_{\mathrm{id}_{M^{\prime}}}$ along $N$ using $\mathrm{id}_{N}$, and then attach the upside down cobordism $\overline{P}$ of $P$ along $M^{\prime}$ using $\mathrm{id}_{M^{\prime}}$. We denote the resulting 4-manifold $Y\cup_{\mathrm{id}_{M}}P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ by $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$. Since the double $\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}$ admits a diffeomorphism to $I\times M^{\prime}$ that fixes the boundary $\overline{M^{\prime}}\sqcup M^{\prime}$ pointwise (see Remark 3.12), the submanifold $P\cup_{\mathrm{id}_{M^{\prime}}}\mathbb{P}\cup_{\mathrm{id}_{N}}\overline{\mathbb{P}}\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ of $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ admits a diffeomorphism to the double $P\cup_{\mathrm{id}_{M^{\prime}}}\overline{P}$ that fixes the boundary $\overline{M}\sqcup M$ pointwise. Let $\widehat{D}_{C_{0}}$ denote the preimage of $D_{C_{0}}$ under this diffeomorphism. By Lemma 10.3, the 4-manifold $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ is diffeomorphic to $Y$ and hence to the exterior $E(X)$, and this diffeomorphism $\varphi_{\mathrm{id}_{M^{\prime}}}:\widehat{Y}_{\mathrm{id}_{M^{\prime}}}\to E(X)$ fixes the boundary $M$ pointwise. We set $D_{C_{0}}^{0}=\varphi_{\mathrm{id}_{M^{\prime}}}(\widehat{D}_{C_{0}})$.
[…]
@@
In the rest, we will prove that the submanifolds $X_{\widetilde{f}}$ are pairwise topologically ambient isotopic in $Z$ by showing that each $X_{\widetilde{f}}$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. Due to the assumption, each $\widetilde{f}$} extends to a self-homeomorphism of $C$ that is identity on the exterior of the contractible 4-manifold $C_{0}$. Extending this homeomorphism as the identity on the rest of $\widehat{Y}_{\widetilde{f}}$, we get a homeomorphism $\widehat{Y}_{\widetilde{f}}\to\widehat{Y}_{\mathrm{id}_{M^{\prime}}}$ that maps $S_{\widetilde{f}}$ to $S_{\mathrm{id}_{M^{\prime}}}$. Note that this homeomorphism is identity on the exterior $\widehat{Y}_{\mathrm{id}_{M^{\prime}}}-\textnormal{int}\,\widehat{D}_{C_{0}}$ of $\widehat{D}_{C_{0}}$. Composing this homeomorphism with the diffeomorphisms $\varphi_{\widetilde{f}}^{-1}$ and $\varphi_{\mathrm{id}_{M^{\prime}}}$, we obtain a self-homeomorphism of $E(X)$ that maps $S^{0}_{\widetilde{f}}$ to $S^{0}_{{\mathrm{id}_{M^{\prime}}}}$ and fixes $E(X)-\textnormal{int}\,D_{C_{0}}^{0}$ and hence the boundary $M$ of $E(X)$ pointwise. Extending this self-homeomorphism as the identity on the rest of $Z$, we obtain a self-homeomorphism ${\Phi}_{\widetilde{f}}$ of $Z$ that maps $X_{\widetilde{f}}$ to $X_{{\mathrm{id}_{M^{\prime}}}}$ and fixes the exterior of $D_{C_{0}}^{0}$ pointwise. Note that ${\Phi}_{\widetilde{f}}$ fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty.
@@
Clearly, the restriction of ${\Phi}_{\widetilde{f}}$ to $D_{C_{0}}^{0}$ fixes $\partial D_{C_{0}}^{0}$ pointwise. Since $D_{C_{0}}^{0}$ is diffeomorphic to $D^{4}$, the Alexander trick [13] shows that the restriction to $D_{C_{0}}^{0}$ is topologically isotopic to the identity map on $D_{C_{0}}^{0}$ fixing $\partial D_{C_{0}}^{0}$ pointwise. Thus, ${\Phi}_{\widetilde{f}}$ is topologically isotopic to the identity map on $Z$. Note that this isotopy fixes $\partial Z$ pointwise in the case where $\partial Z$ is non-empty. Therefore, each submanifold $X_{\widetilde{f}}$ of $Z$ is topologically ambient isotopic to $X_{{\mathrm{id}_{M^{\prime}}}}$. This completes the proof. ∎
@@
Let $Y$ be a 4-manifold diffeomorphic to the exterior $E(X)$. We first consider the case where $Y$ admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Let $(C_{0},\tau)$ be an infinite order contractible cork of Gompf (see Subsection 3.5). By results of Akbulut [5], Tange [66] and Naylor and Schwartz [57], the twisted double $C_{0}\cup_{\tau^{i}}\overline{C_{0}}$ is diffeomorphic to $S^{4}$ for each integer $i$. Therefore, we can apply Proposition 10.4 using a generalized cork obtained from $Y$ in the proof of Proposition 4.2. The claim thus follows in this case.
We next consider the case where $Y$ is a 2-handlebody. In this case, we modify $Y$ into a Stein handlebody $Y^{\prime}$ by using Proposition 3.6. Note that $Y$ is obtained from $Y^{\prime}$ by attaching the upside down cobordism of an invertible cobordism $P$. Furthermore, $P$ admits a deformation retraction onto the negative boundary $\partial Y$. Since $Y$ is diffeomorphic to $E(X)$, we may regard $P$ as a submanifold of $E(X)$ such that the negative boundary of $P$ is $\partial E(X)$. Then, the union $X^{\prime}=X\cup P$ is a submanifold of $Z$. Clearly, $X^{\prime}$ contains $X$ as a submanifold and admits a deformation retraction onto $X$. Since the exterior of $X^{\prime}$ in $Z$ is diffeomorphic to the Stein manifold $Y^{\prime}$, we can prove the desired claim by applying that of the first case to $X^{\prime}$. ∎
@@
For Theorem 10.2 and Proposition 10.4, the exotically knotted submanifolds of $Z$ are obtained from $X$ by attaching invertible homology cobordisms embedded in $Z$, and hence they admit embeddings into $X$. It thus follows from Corollary 7.5 that the genus functions of these exotically knotted submanifolds are algebraically equivalent to that of $X$.
@@
Let $Z$ be a (possibly non-orientable) 4-manifold which may have (possibly disconnected) boundary, and let $M_{1},M_{2}$ be codimension-1 submanifolds of $Z$. The submanifold $M_{1}$ is called homology cobordant to $M_{2}$ through $Z$ if $Z$ admits a codimension-0 (oriented) submanifold $P$ with $\partial P=\overline{M_{1}}\sqcup M_{2}$ such that $P$ is a homology cobordism from $M_{1}$ to $M_{2}$.
As an application of Theorem 10.2, we modify separating codimension-$1$ submanifolds into infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds. Specifically, we prove the following theorem which is a refinement of Theorem 1.14.
@@
Let $M$ be a closed $3$-manifold embedded in the interior of a ((possibly non-orientable)) 44-manifold $Z$ which may have ((possibly disconnected)) boundary. Suppose the exterior of the tubular neighborhood of $M$ has a connected component with connected boundary which either is a $4$-dimensional ((orientable)) 22-handlebody or admits an embedding into a symplectic $4$-manifold with weakly convex boundary. Then, $Z$ admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$ through $Z$.
@@
Due to the assumption, we can apply Theorem 10.2 to a component $X$ of the exterior of the tubular neighborhood for $M$. Note that $\partial E(X)$ is diffeomorphic to $M$. Let $X_{i}$ $(i\in\mathbb{Z})$ be the resulting infinitely many pairwise exotically knotted and exotically embedded codimension-0 (possibly non-orientable) submanifolds of $Z$, and let each $M_{i}$ be the boundary $\partial E(X_{i})$ of the exterior $E(X_{i})$.
We fix an orientation of the exterior $E(X)$ and equip each $E(X_{i})\subset E(X)$ with the orientation induced from $E(X)$. In the case where the orientaion of $M$ is that of $\partial E(X)$ (resp. $\partial\overline{E(X)}$), we equip each $M_{i}$ with the orientaion of $\partial E(X_{i})$ (resp. $\partial\overline{E(X_{i})}$). Clearly, the 3-manifolds $M_{i}$ are pairwise orientedly diffeomorphic. By the proof of Theorem 10.2, each $M_{i}$ is homology cobordant to $M$ through $Z$. Since the submanifolds $X_{i}$ are pairwise topologically ambient isotopic, their boundary components $M_{i}$ are also pairwise topologically ambient isotopic.
We will show that the pairs $(Z,M_{i})$ are pairwise non-diffeomorphic. By the proof of Theorem 10.2, the exteriors $E(X_{i})$ are pairwise exotic, and the 4-manifolds $X_{i}$ are pairwise diffeomorphic. Note that the exterior of the tubular neighborhood of each $M_{i}$ is the disconnected 4-manifold $E(X_{i})\sqcup X_{i}$. Now suppose to the contrary that a pair $(Z,M_{i})$ is diffeomorphic to some $(Z,M_{j})$ with $i\neq j$. Then, $E(X_{i})\sqcup X_{i}$ is diffeomorphic to $E(X_{j})\sqcup X_{j}$ due to the uniqueness of the tubular neighborhood. Since $E(X_{i})$ is not diffeomorphic to $E(X_{j})$, we see that $E(X_{i})$ is diffeomorphic to $X_{j}$ and hence to $X_{i}$ and that $X_{i}$ is diffeomorphic to $E(X_{j})$. Consequently, $E(X_{i})$ is diffeomorphic to $E(X_{j})$, giving a contradiction. Thus, the pairs $(Z,M_{i})$ are pairwise non-diffeomorphic. Therefore, the 3-manifolds $M_{i}$ are pairwise exotically embedded and exotically knotted in $Z$. ∎
We obtain the following corollary which is a restatement of Corollary 1.15.
@@
For any closed $3$-manifold $M$ embedded in $S^{4}$, every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded codimension-$1$ submanifolds which are homology cobordant to $M$.
@@
Let $M$ be a closed 3-manifold embedded in $S^{4}$. It follows from the Mayer–Vietoris exact sequence that $M$ is separating. Let $Y$ be a connected component of the exterior of the tubular neighborhood for $M$. Since $M$ is separating, the boundary $\partial Y$ is diffeomorphic to $M$. Also, $Y$ admits an embedding into $D^{4}$ and hence into a symplectic 4-manifold with weakly convex boundary. Now let $Z$ be an arbitrary (possibly non-orientable) 4-manifold. Then $Y$ admits an embedding into the interior of $Z$, and thus we can apply Theorem 10.7 to the submanifold $\partial Y$ of $Z$. Therefore, the claim follows. ∎
The above corollary implies the following corollary which is a refinement of Corollary 1.16.
@@
For each integer $n\geq 0$, there exists a closed $3$-manifold $M$ with $b_{1}=n$ such that every ((possibly non-orientable)) 44-manifold admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$.
@@
As is well-known, for each integer $n\geq 0$, there are various closed 3-manifolds with $b_{1}=n$ admitting embeddings into $S^{4}$. For example, each $n(S^{1}\times S^{2})$ is such a 3-manifold. The claim thus follows from Corollary 10.8. ∎
For a given closed 3-manifold, here we consider the question of what is the minimum second Betti number of a closed 4-manifold which contains a pair of exotically knotted (or exotically embedded) separating submanifolds diffeomorphic to the 3-manifold. Since every closed 3-manifold embedded in a simply connected closed 4-manifold is separating as follows from the Mayer–Vietoris exact sequence, this question is an extension of the question mentioned in the paragraph preceding Corollary 1.17, where the ambient 4-manifold is assumed to be simply connected. As an application of Theorem 10.7, we show that every non-negative integer is realized as such a minimum number, giving a refinement of Corollary 1.17.
@@
For each even integer $n\geq 0$ and each integer $m$ with $m\geq n/2$, there exists a closed $3$-manifold $M$ with $b_{1}=m$ such that a ((simply connected)) closed $4$-manifold with $b_{2}=n$ contains infinitely many pairwise exotically knotted and exotically embedded separating submanifolds diffeomorphic to $M$, but no closed $4$-manifold with $b_{2}<n$ contains a separating submanifold diffeomorphic to $M$.